40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Mon Jul 13 13:59:19 2009, Michelle, Laser, General, NPRO acting up 
    Reply  Mon Jul 13 21:39:11 2009, Michelle, Laser, General, NPRO acting up 
    Reply  Wed Jul 15 10:47:55 2009, Michelle, Laser, General, NPRO acting up 
Message ID: 172     Entry time: Wed Jul 15 10:47:55 2009     In reply to: 169
Author: Michelle 
Type: Laser 
Category: General 
Subject: NPRO acting up 


I'm having some issues with the 495 mW NPRO, or possibly with its power supply. I ran the laser for about 15 minutes at 2 mW while doing some aligning. I closed the shutter to move a mirror, and upon opening the shutter the laser was no longer lasing. The power supply read 0.0 A driving current. I closed the shutter again and checked all connections between the laser and the power supply. All in order. I opened the shutter and turned the dial a bit on the power supply, and lo and behold the current was back up to 0.46 A, the lasing threshold, but still no light coming out.

The drive current jumped around a lot, sometimes to 0.0 A, and even when I put it up to around 0.54 A, which the laser should definitely be lasing at, nothing came out the aperture. The internal laser temperature was normal and not fluctuating. The lab itself is pretty warm today, in the 70s. Unless someone changed the temperature, it felt like the heat was on. I don't think this would have any effect on the laser, but I'm not really sure.

So, I'm going to search for the manual online and see if anyone over at the 40m had problems with this power supply.

 I took the power supply over to the 40m to see if Alberto's laser would work when hooked up to it. It did, so I suspect that the problem lies with the laser rather than the power supply. Alberto was able to get double the drive current that we were getting, so that is also suspect. He said the "ADJ" readout (power adjustment) on the supply was usually set to zero, but when the "Pwr" readout (supposedly the power coming out of the laser) is changed, it changes the ADJ and drive current and anything else which would obviously be associated with power. I hooked up the power supply to my laser again and here's what I found:

-It lases for short periods of time, usually <5 mins, before it stops.

-It lases for longer periods of time at higher drive currents.

-I still can't get more than 1.02 A and less than 0.30 A out of the power supply.

-The "Pwr" display on the supply still doesn't correspond to the power as measured by the Ophir powermeter.

My next steps are to see if Alberto can come and take a look, in case I'm doing something on the operator side which is obviously wrong. I will also consult with Rana.

ELOG V3.1.3-